# 2.5cm2.5cm2cm2cm

### v1

@@ 3.2. The Nash discriminant variety of a three-player binary game
$i$ $\mathrm{Sing}^{(i)}\Delta$ $\dim\mathrm{Sing}^{(i)}\Delta${irreducible components}
$1$ $\Delta^{\mathrm{cub}}\cup\Delta^{\mathrm{con}}\cup\Delta^{\mathrm{lin}}$ $8$ $7=1+3+3$
$2$ $\Delta^{\mathrm{ll}}\cup\Delta^{\mathrm{cl}}$ $7$ $6=3+3$
$3$ $\Delta^{\mathrm{lll}}\cup\Delta^{\mathrm{scr}}$ $6$ $6=3+3$
$4$ $\Delta^{\mathrm{3l}}$ $5$ $1$
$5$ $\Delta^{\mathrm{qq}}$ $2$ $3$
@@
[…]
@@
$X^{(1)}=\begin{bmatrix}[cc:cc]-8&12&10&-15\\ 50&0&0&0\end{bmatrix}\,,\ X^{(2)}=\begin{bmatrix}[cc:cc]-2&0&3&0\\ 2&0&-3&0\end{bmatrix}\,,\ X^{(3)}=\begin{bmatrix}[cc:cc]-5&5&0&0\\ -5&10&0&0\end{bmatrix}\,,$
@@ 3.3. The Nash discriminant variety of games of boundary format
This subsection concerns the games of boundary format (the games at the boundary between the balanced and unbalanced games). This subsection aims to study the Nash discriminant variety of such games and to reveal its unexpected properties. To be more specific, we prove that the Nash discriminant variety is reducible. One of the irreducible components is a hypersurface. We also discuss a formula for the degree of this hypersurface component.
@@
Next, suppose that $[f]\in\Delta(E)\setminus{\mathcal{D}}_{1}(E)$. Let $f=(f^{(1)},\ldots,f^{(n)})$ be the decomposition of $f$ into global sections of ${\mathcal{O}}({\boldsymbol{1}}_{i})^{\oplus(d_{i}-1)}$. Consider $f^{(n)}$ as a global section of $E^{\prime}$. The zero scheme ${\mathcal{Z}}(f^{(n)})$ is nonsingular of codimension $d_{n}-1$, because $E^{\prime}$ is globally generated and $[f]\not\in{\mathcal{D}}_{1}$ (and hence $[f^{(n)}]\not\in\Delta(E^{\prime})$). Furthermore, it consists of $c_{d_{n}-1}(E^{\prime})=(d_{n}-1)!/((d_{1}-1)!\cdots(d_{n-1}-1)!)$ distinct closed points.
The canonical projection from ${\mathbb{P}}^{\boldsymbol{d}}$ to ${\mathbb{P}}^{{\boldsymbol{d}}^{\prime}}$ maps ${\mathcal{Z}}(f)$ onto ${\mathcal{Z}}(f^{(n)})$. The fiber of the projection restricted to ${\mathcal{Z}}(f)$ over a closed point of ${\mathcal{Z}}(f^{(n)})$ is the linear subspace of ${\mathbb{P}}^{d_{n}-1}$ defined by the $d_{n}-1$ linear forms in $\pi^{(n)}$ obtained from $f^{(1)},\ldots,f^{(n-1)}$ by evaluating at the closed point. Therefore, the scheme ${\mathcal{Z}}(f)$ is the disjoint union of $(d_{n}-1)!/((d_{1}-1)!\cdots(d_{n-1}-1)!)$ (possibly different dimensional) linear subspaces. This means that $[f]\in\Delta(E)\setminus{\mathcal{D}}_{1}$ if and only if ${\mathcal{Z}}(f)$ has a positive dimensional component.
@@
As was shown above, if $[f]\in\Delta(E)\setminus{\mathcal{D}}_{1}$, then $\dim{\mathcal{Z}}(f)\geq 1$. Therefore,
@@
we obtain the generators of the ideal of the second component ${\mathcal{D}}_{2}$ of $\Delta(E)$. With the aid of Macaulay2, one can verify that ${\rm codim}\,{\mathcal{D}}_{2}=2$ and $\deg{\mathcal{D}}_{2}=19$. One can also check that ${\mathcal{D}}_{2}$ is irreducible over the field ${\mathbb{Q}}$ and its radical ideal is minimally generated by $11$ homogeneous polynomials in $a_{ij}^{(1)},a_{ij}^{(2)},a_{ij}^{(3,1)}$, and $a_{ij}^{(3,2)}$; three of them have degree $6$, two of them have degree $7$, and the rest of them have degree $8$diamondsuit
@@
is a solution to the system mentioned above. With the aid of Macaulay2, we verified that ${\mathcal{Z}}_{X}$ is a nonreduced point of multiplicity $2$ supported at $[{\boldsymbol{\pi}}]\in{\mathbb{P}}^{\boldsymbol{d}}$. $\diamondsuit$
@@ 3.4. The Nash resultant variety
Let ${\boldsymbol{d}}=(d_{1},d_{2})\in{\mathbb{Z}}_{{\geq 2}}^{2}$ with $d_{1}<d_{2}$. The Nash resultant variety ${\mathcal{R}}({\boldsymbol{d}})$ is irreducible of codimension $d_{2}-d_{1}+1$ and degree $\binom{d_{2}-1}{d_{1}-1}$.
@@
Since $d_{1}<d_{2}$, this system is overdetermined, and hence ${\mathcal{Z}}_{X}\neq\emptyset$ precisely when $\mathrm{rank}\,A\leq d_{1}-1$. In particular, the Nash resultant variety ${\mathcal{R}}({\boldsymbol{d}})$ is a determinantal variety, and its codimension and degree are ${\rm codim}\,{\mathcal{R}}({\boldsymbol{d}})=d_{2}-d_{1}+1$ and $\deg{\mathcal{R}}({\boldsymbol{d}})=\binom{d_{2}-1}{d_{1}-1}$ respectively. ∎
@@
$\deg{\mathcal{R}}({\boldsymbol{d}})$ $=\binom{d_{n}-1}{{\rm codim}\,{\mathcal{R}}({\boldsymbol{d}})-1}\binom{d_{n}-{\rm codim}\,{\mathcal{R}}({\boldsymbol{d}})-1}{d_{1}-1,\ldots,d_{n-1}-1}$
@@
it suffices to show that ${\mathcal{R}}(F)$ is irreducible and to compute its codimension and degree.
If $d_{n}-1=\sum_{i=1}^{n-1}(d_{i}-1)+1$, then it follows from [GKZ94, Chapter 3, Proposition 3.1] that ${\mathcal{R}}(F)$ is an irreducible hypersurface in $H^{0}({\mathbb{P}}^{{\boldsymbol{d}}^{\prime}},F)$. Furthermore, since $F$ is very ample,
$\deg{\mathcal{R}}(F)=\int_{{\mathbb{P}}^{{\boldsymbol{d}}^{\prime}}}c_{d_{n}-2}(F)$
by [GKZ94, Chapter 3, Theorem 3.10]. Therefore, $\deg{\mathcal{R}}(F)$ is equal to the coefficient of $\prod_{i=1}^{n-1}h_{i}^{d_{i}-1}$ in the polynomial $\hat{h}_{n}^{d_{n}-2}$, where $\hat{h}_{j}=\sum_{i\neq j}h_{i}$. In particular,
$\deg{\mathcal{R}}(F)$ $=\binom{d_{n}-1}{d_{n}-2}\binom{d_{n}-2}{d_{1}-1,d_{2}-1,\ldots,d_{n-1}-1}$
$=\frac{(d_{n}-1)!}{(d_{1}-1)!(d_{2}-1)!\cdots(d_{n-1}-1)!}\,.$
With slight modifications, the proofs of [GKZ94, Chapter 3, Proposition 3.1, Theorem 3.10] show that, if $d_{n}-1>\sum_{i=1}^{n-1}(d_{i}-1)+1$, then ${\mathcal{R}}(F)$ is irreducible of codimension $c\coloneqq d_{n}-1-\sum_{i=1}^{n-1}(d_{i}-1)$ and that its degree is
@@ Acknowledgements
This work is partially supported by the Thematic Research Programme “Tensors: geometry, complexity and quantum entanglement”, University of Warsaw, Excellence Initiative - Research University and the Simons Foundation Award No. 663281 granted to the Institute of Mathematics of the Polish Academy of Sciences for the years 2021-2023. We are thankful for the support and great working conditions during the semester “Algebraic Geometry with Applications to Tensors and Secants” (AGATES). HA would like to express gratitude to the Max Planck Institute for Mathematics in the Sciences for its generous hospitality and inspiring research environment during his visit. IP expresses her sincere gratitude to KTH Royal Institute of Technology in Stockholm for its hospitality and stimulating discussions during her visit. The work of LS was partially supported by a KTH grant from the Verg Foundation and Brummer & Partners MathDataLab. We are grateful to Antonio Lerario for fruitful discussions on Thom’s Isotopy Lemma.

### v2

@@ 3.2. The Nash discriminant variety of a three-player binary game
$i$ $\mathrm{Sing}^{(i)}\Delta$ $\dim\mathrm{Sing}^{(i)}\Delta$ #{irreducible components}
11 $\Delta^{\mathrm{cub}}\cup\Delta^{\mathrm{con}}\cup\Delta^{\mathrm{lin}}$ 88 $7=1+3+3$
22 $\Delta^{\mathrm{ll}}\cup\Delta^{\mathrm{cl}}$ 77 $6=3+3$
33 $\Delta^{\mathrm{lll}}\cup\Delta^{\mathrm{scr}}$ 66 $6=3+3$
44 $\Delta^{\mathrm{3l}}$ 55 11
55 $\Delta^{\mathrm{qq}}$ 22 33
@@
If $\pi^{(1)}=\left(\gamma_{2}/(\gamma_{2}-\gamma_{1}),-\gamma_{1}/(\gamma_{2}-\gamma_{1})\right)\in\Delta_{1}^{\circ}$, if $\pi^{(2)}\in\Delta_{1}^{\circ}$ satisfies $\alpha_{1}\beta_{1}+\alpha_{2}\beta_{2}\penalty 10000\ =\penalty 10000\ 0$, and if
@@
$X^{(1)}=\begin{bmatrix}[cc:cc]-8&12&10&-15\\ 0&0&0&0\end{bmatrix}\,,\ X^{(2)}=\begin{bmatrix}[cc:cc]-2&0&3&0\\ 2&0&-3&0\end{bmatrix}\,,\ X^{(3)}=\begin{bmatrix}[cc:cc]-5&5&0&0\\ -5&10&0&0\end{bmatrix}\,$,
@@ 3.3. The Nash discriminant variety of games of boundary format
This subsection concerns games of boundary format (games at the boundary between the balanced and unbalanced games). This subsection aims to study the Nash discriminant variety of such games and to reveal its unexpected properties. To be more specific, we prove that the Nash discriminant variety is reducible. One of the irreducible components is a hypersurface. We also discuss a formula for the degree of this hypersurface component.
@@
Next, suppose that $f]\in\Delta(E)\setminus{\mathcal{D}}_{1}(E)$. Let $f=(f^{(1)},\ldots,f^{(n)})$ be the decomposition of $f$ into global sections of ${\mathcal{O}}({\boldsymbol{1}}_{i})^{\oplus(d_{i}-1)}$. Consider $f^{(n)}$ as a global section of $E^{\prime}$. The zero scheme ${\mathcal{Z}}(f^{(n)})$ is nonsingular of codimension $d_{n}-1$, because $E^{\prime}$ is globally generated and $f]\not\in{\mathcal{D}}_{1}$ (and hence $f^{(n)}]\not\in\Delta(E^{\prime})$). Furthermore, it consists of $c_{d_{n}-1}(E^{\prime})=(d_{n}-1)!/((d_{1}-1)!\cdots(d_{n-1}-1)!)$ distinct closed points.
The canonical projection from ${\mathbb{P}}^{\boldsymbol{d}}$ to ${\mathbb{P}}^{{\boldsymbol{d}}^{\prime}}$ maps ${\mathcal{Z}}(f)$ onto ${\mathcal{Z}}(f^{(n)})$. The fiber of the projection restricted to ${\mathcal{Z}}(f)$ over a closed point of ${\mathcal{Z}}(f^{(n)})$ is the linear subspace of ${\mathbb{P}}^{d_{n}-1}$ defined by the $d_{n}-1$ linear forms in $\pi^{(n)}$ obtained from $f^{(1)},\ldots,f^{(n-1)}$ by evaluating at the closed point. Therefore, the scheme ${\mathcal{Z}}(f)$ is the disjoint union of $(d_{n}-1)!/((d_{1}-1)!\cdots(d_{n-1}-1)!)$ (possibly different dimensional) linear subspaces. This means that $f]\in\Delta(E)\setminus{\mathcal{D}}_{1}$ if and only if ${\mathcal{Z}}(f)$ has a positive dimensional component.
@@
As was shown above, if $f]\in\Delta(E)\setminus{\mathcal{D}}_{1}$, then $\dim{\mathcal{Z}}(f)\geq 1$. Therefore,
@@
we obtain the generators of the ideal of the second component ${\mathcal{D}}_{2}$ of $\Delta(E)$. With the aid of Macaulay2, one can verify that ${\rm codim}\,{\mathcal{D}}_{2}=2$ and $\deg{\mathcal{D}}_{2}=19$. One can also check that ${\mathcal{D}}_{2}$ is irreducible over the field $\mathbb{Q}$} and its radical ideal is minimally generated by $11$ homogeneous polynomials in $a_{ij}^{(1)},a_{ij}^{(2)},a_{ij}^{(3,1)}$, and $a_{ij}^{(3,2)}$; three of them have degree $6$, two of them have degree $7$, and the rest of them have degree $8$. $\diamondsuit$
@@
is a solution to the system mentioned above. With the aid of Macaulay2, we verified that ${\mathcal{Z}}_{X}$ is a nonreduced point of multiplicity $2$ supported at ${\boldsymbol{\pi}}]\in{\mathbb{P}}^{\boldsymbol{d}}$. $\diamondsuit$
@@ 3.4. The Nash resultant variety
Let ${\boldsymbol{d}}=(d_{1},d_{2})\in{\mathbb{Z}}_{{\geq 2}}^{2}$ with $d_{1}<d_{2}$. The Nash resultant variety ${\mathcal{R}}({\boldsymbol{d}})$ is irreducible of codimension $d_{2}-d_{1}$ and degree $\binom{d_{2}-1}{d_{1}-1}$.
@@
Since $d_{1}<d_{2}$, this system is overdetermined, and hence ${\mathcal{Z}}_{X}\neq\emptyset$ precisely when $\mathrm{rank}\,A\leq d_{1}-1$. In particular, the Nash resultant variety ${\mathcal{R}}({\boldsymbol{d}})$ is a determinantal variety, and its codimension and degree are ${\rm codim}\,{\mathcal{R}}({\boldsymbol{d}})=d_{2}-d_{1}$ and $\deg{\mathcal{R}}({\boldsymbol{d}})=\binom{d_{2}-1}{d_{1}-1}$ respectively. ∎
@@
$\deg{\mathcal{R}}({\boldsymbol{d}})$ $=\binom{d_{n}-1}{{\rm codim}\,{\mathcal{R}}({\boldsymbol{d}})}\binom{d_{n}-{\rm codim}\,{\mathcal{R}}({\boldsymbol{d}})-1}{d_{1}-1,\ldots,d_{n-1}-1}$
@@
it suffices to show that ${\mathcal{R}}(F)$ is irreducible and to compute its codimension and degree. These were already studied in [GKZ94, Section 3 in Chapter 3]. For the sake of the reader’s convenience, we outline the proofs used in [GKZ94] below.
Let $\Gamma(F)\coloneqq\{(p,[f^{(n)}])\in{\mathbb{P}}^{{\boldsymbol{d}}^{\prime}}\times{\mathbb{P}}H^{0}({\mathbb{P}}^{{\boldsymbol{d}}^{\prime}},F)\mid{\mathcal{Z}}(f^{(n)})\neq\emptyset\}$, and for each $i\in\{1,2\}$, denote the projection from ${\mathbb{P}}^{{\boldsymbol{d}}^{\prime}}\times{\mathbb{P}}H^{0}({\mathbb{P}}^{{\boldsymbol{d}}^{\prime}},F)$ to its $i$th factor by $q_{i}$. The fiber $q_{1}^{-1}(p)$ of $q_{1}$ over any point $p$ of ${\mathbb{P}}^{{\boldsymbol{d}}^{\prime}}$ is identified with a linear subspace of ${\mathbb{P}}H^{0}({\mathbb{P}}^{{\boldsymbol{d}}^{\prime}},F)$ defined by $d_{n}-1$ linearly independent linear forms obtained by evaluating $f^{(n)}$ at $p$. This means that $q_{1}^{-1}(p)$ has codimension $d_{n}-1$, and hence $\Gamma(F)$ has a projective bundle structure with the morphism $q_{1}\colon\Gamma(E)\to{\mathbb{P}}^{{\boldsymbol{d}}^{\prime}}$. In particular, it is irreducible and has $\dim{\mathbb{P}}^{{\boldsymbol{d}}^{\prime}}+\dim{\mathbb{P}}H^{0}({\mathbb{P}}^{{\boldsymbol{d}}^{\prime}},F)-(d_{n}-1)=\dim{\mathbb{P}}H^{0}({\mathbb{P}}^{{\boldsymbol{d}}^{\prime}},F)-(d_{n}-1-\sum_{i=1}^{n-1}(d_{i}-1))$. Furthermore, since $F$ is very ample, if $f^{(n)}]\in{\mathcal{R}}(F)$ is generic, then $Z(f^{(n)})$ consists of a closed point. Therefore, the mprphism $q_{2}\colon\Gamma(F)\to{\mathcal{R}}(F)$ is birational, and hence ${\mathcal{R}}(F)$ is irreducible of codimension $d_{n}-1-\sum_{i=1}^{n-1}(d_{i}-1)$.
Regarding the degree of ${\mathcal{R}}(F)$, the proof of [GKZ94, Chapter 3, Theorem 3.10] shows that, if $d_{n}-1\geq\sum_{i=1}^{n-1}(d_{i}-1)+1$, then its degree is
@@ Acknowledgements
This work is partially supported by the Thematic Research Programme “Tensors: geometry, complexity and quantum entanglement”, University of Warsaw, Excellence Initiative - Research University and the Simons Foundation Award No. 663281 granted to the Institute of Mathematics of the Polish Academy of Sciences for the years 2021-2023. We are thankful for the support and excellent working conditions during the semester “Algebraic Geometry with Applications to Tensors and Secants” (AGATES). HA would like to express gratitude to the Max Planck Institute for Mathematics in the Sciences for its generous hospitality and inspiring research environment during his visit. IP expresses her sincere gratitude to KTH Royal Institute of Technology in Stockholm for its hospitality and stimulating discussions during her visit. The work of LS was partially supported by a KTH grant from the Verg Foundation and Brummer & Partners MathDataLab. We are grateful to Antonio Lerario for fruitful discussions on Thom’s Isotopy Lemma. The authors also thank the referee and Claus Hertling for carefully reading the paper and for their valuable suggestions and comments.